# Gordian split links in the Gehring ropelength problem

### v1

@@ Abstract
A thick link is a link in $\mathbb{R}^{3}$ such that each component of the link lies at distance at least $1$ from every other component. Strengthening the notion of thickness, we define a thickly embedded link to be a thick link whose open radius-$\tfrac{1}{2}$ normal disk bundles of all components are embedded. The Gehring ropelength problem asks how large the sum of the lengths of the components of a thick (respectively thickly embedded) link must be, given the link homotopy (respectively isotopy) class of the link. A thick homotopy (isotopy) is a link homotopy (isotopy) of a thick (thickly embedded) link that preserves thickness throughout, and such that during the homotopy the total length of the link never exceeds the initial total length. These notions of thick homotopy and isotopy are more permissive than other notions of physical link isotopies in which the length of each individual component must remain constant (no “length trading”). We construct an explicit example of a thickly embedded 4-component link which is topologically split but cannot be split by a thick homotopy, and thick links in every homotopy class with 2 components that are non-global local minima for ropelength. This is the first time such local minima for ropelength have been explicitly constructed. In particular, we construct a thick 2-component link in the link homotopy class of the unlink which cannot be split through a thick homotopy.
[…]
[inst1]organization=Dartmouth College Department of Mathematics
@@ 1 Introduction
In 2015, Coward and Hass [5] defined the notion of a physical link isotopy. A physical link isotopy is a link isotopy that keeps a link thickly embedded throughout, as well as keeping the length of each component constant. This made it possible to study a physical knot- and link theory that corresponds to knots and links tied from real pieces of rope. In the same paper, Coward and Hass construct a $2$-component Gordian split link which they define as a thickly embedded, topologically split link that cannot be split by a physical link isotopy. More generally, a Gordian pair of links is a set of two links in $\mathbb{R}^{3}$ which are ”topologically equivalent”, but which are not equivalent under ”physical deformations”. Here ”topologically equivalent” refers either to being in the same link isotopy or homotopy class, and ”physical deformation” refers to either link homotopy or link isotopy, subject to additional geometric constraints on length and thickness. For length, there are two common constraints: either demanding that the length of each component individually does not increase throughout the homotopy/isotopy, or that their sum does not increase throughout the homotopy/isotopy. If only the sum of lengths is constrained, we say that ”length trading” between components is permitted. The thickness constraint usually demands that links stay thick links (see Definition 3.3) or thickly embedded links (see Definition 3.4) throughout the homotopy or isotopy respectively. When considering link isotopy, sometimes additional curvature constraints are imposed.
@@
There are good reasons to study thick homotopies even though they are the least physical of the ”physical deformations”. First, once it is proven that a pair of ”topologically equivalent” links does not have a thick homotopy between them, the pair of links will a fortiori be a Gordian pair in all other reasonable notions of physicality. Secondly, the lack of thick homotopies between two links in the same link homotopy class is related to local minima in the general Gehring ropelength problem: Two links in a given link homotopy class, one minimizing ropelength globally and one only locally, should not have a thick homotopy between them. This will be the strategy we use to construct Gordian pairs in every link homotopy class of two components.
@@ 2 Overview of existing results about Gordian pairs
Any Gordian pair Split link $2$-component split link $2$-component unlink
@@
Any Gordian pair Split link $2$-component unlink All $2$-component links
@@
The first Gordian unlink was constructed by Coward and Hass in 2015 [5]. In 2023, Kusner and Kusner [8] constructed a Gordian pair of thickly embedded $7$-component links. The links in the pair are link isotopic to each other, but there is no thick homotopy between them. So even though the links are equivalent in the strongest topological sense, they are not equivalent even in the weakest ”physical” sense. It was Kusner and Kusner who first coined the term Gordian pair. In [1] Ayala constructs a thickly embedded $2$-component link in the isotopy class of the unlink which cannot be split by thick isotopy under additional curvature constraints. This was the first example of a Gordian unlink. Recently, Ayala and Hass constructed a different example of a Gordian unlink without needing any additional curvature constraints. Independently, Kusner and Kusner announced the construction of a Gordian unlink [9]. Table 1 summarizes both the existing results on Gordian pairs and the author’s contributions in this paper. The table is split into two parts, considering links with respect to isotopy and homotopy respectively. The columns of each table denote what kind of example was constructed. The rows of each table denote which geometric constraints were considered, with the strength of constraints being relaxed from top to bottom. The paper referenced in each cell is the paper in which the first example of the corresponding Gordian pair was constructed. The tables give partial orders: the further down and to the right a cell in a given table is, the harder it is to construct an example for it.
@@ 3 Preliminaries
Let $A,B\subseteq\mathbb{R}^{n}$ be non-empty sets, let $x\in\mathbb{R}^{n}$ and let $f,g$ be functions with image in $\mathbb{R}^{n}$. Then we denote the distance of $A$ to $B$ with $\mathrm{dist}(A,B)\coloneqq\inf\{||a-b||\mid a\in A,b\in B\}$; and we define $\mathrm{dist}(f,g)$, $\mathrm{dist}(f,A)$, $\mathrm{dist}(x,A)$, $\mathrm{dist}(x,f)$ by identifying $x$ with $\{x\}$ and $f,g$ with their image. If $f,g:M\to\mathbb{R}^{n}$ then $||f-g||\coloneqq\sup\{||f(x)-g(x)||\mid x\in M\}$.
@@
$||L-L^{\prime}||<\varepsilon\implies\mathrm{len}(L^{\prime})\geq\mathrm{len}(L).$
@@
If a thick link $L$ is a minimum of the preorder $\preceq$, then we call $L$ a sink of the ropelength functional.
@@
$(f\circ\gamma)(t)\leq f(x)\text{ for all t}\implies f\circ\gamma\text{ constant. }$.
It is easy to see that every global minimum and every strict local minimum is a sink. However, even for ”nice” spaces $X$ and continuous functions $f$, there are local minima which are not sinks and sinks which are not local minima.
@@
and with equality if and only if $\gamma(t^{\prime})\in S$. By assumption $f(\gamma(t^{\prime}))\leq c$. Therefore $f(\gamma(t^{\prime}))=c$ and hence $\gamma(t^{\prime})\in S$ This shows that $J$ is open. Since $J$ is a non-empty, closed and open subset of $[0,1]$ we have $J=[0,1]$. So $\gamma([0,1])\subseteq S\subseteq f^{-1}(c)$, hence $f\circ\gamma$ is constant. ∎
@@ 4 Bounds on ropelength
Let $x^{\prime},y^{\prime}\in K$ with $||x^{\prime}-x||,||y^{\prime}-y||\leq\varepsilon$. Then $\mathrm{dist}(x^{\prime},C),\mathrm{dist}(y^{\prime},C)\geq 1-\varepsilon.$ Hence by Proposition 4.24
$\mathrm{len}(C)\geq$ $2\pi(1-\varepsilon)+2||x^{\prime}-y^{\prime}||$
$\geq$ $2\pi(1-\varepsilon)+2||x-y||-4\varepsilon$
$=$ $2\pi+2||x-y||-(2\pi+4)\varepsilon.$
@@ 5 An isotopically Gordian split link exists
The construction of our Gordian split link, as well as the basic structure of the proof, has a very similar flavor to that of Coward and Hass [5]. Both approaches derive a lower bound for the length needed to split a thickly embedded link by assuming that there is a thick homotopy/physical isotopy, tracking the intersections of a center-of-mass spanning disk and a curve throughout that homotopy/isotopy, and finally arguing that there needs to be some point in time where the intersections yield some lower bound through applications of propositions from section 4. The main difference is in the ”book-keeping” of how intersections are tracked. Coward and Hass need to resolve multiple kinds of singularities that might occur during the tracking of intersections. Our arguments use the homotopy invariance of the Brouwer degree and the mountain climbing theorem instead. We start out by stating the main proposition of this chapter, which establishes a lower-bound for the length necessary to split some thick links.
@@ 5.1 A lower bound on ropelength needed to split some thick links
The main step omitted in this sketch is how, given a homotopy $H$ preserving thickness, we can continuously keep track of the intersection points $a_{t}$ and $b_{t}$, and why we can assume that they merge eventually. In this form the claim is wrong, but we will see that it can be replaced with a more technical, true claim in the same spirit. The first tool we need to use for our book-keeping will be the homotopy invariance of the Brouwer degree of a function around 00. We will therefore briefly recall the following existence and uniqueness theorem defining the degree. This formulation is taken from Fixed Point Theory by Granas and Dugundji [6].
5.29 Theorem (Brouwer degree around 00).
@@
=$\mathcal{M}=\{(U,F)\mid U\subseteq\mathbb{R}^{n}\text{ open, bounded },F:\overline{U}\to\mathbb{R}^{n}\text{ continuous},0\notin F(\partial U)\}.$
@@
We further need the following two facts about the Brouwer degree. If $F:U\to\mathbb{R}^{n}$ is $\mathcal{C}^{1}$ and if 00 is a regular value of $F$, then
@@
If $0\notin F(\partial U)$, then $00\notin F(U)\implies d(U,F)=$ 0 .
@@
Let $D^{2}$ be a disk and let $K,L\subseteq D^{2}$ be closed such that $D^{2}=K\cup L$ and such that $K\cap S^{1}\times\{0\}=I_{A}\cup I_{B}$ and $L\cap S^{1}\times\{0\}=I_{X}\cup I_{Y}$ are each the disjoint union of two compact intervals of positive length, with $I_{A}\cap I_{X}$, $I_{B}\cap I_{X}$, $I_{A}\cap I_{Y}$, and $I_{B}\cap I_{Y}$ each containing exactly one element. Then either $I_{A}$ and $I_{B}$ are in the same connected component of $K$ or $I_{X}$ and $I_{Y}$ are in the same connected component of $L$.
@@
This lemma tells us that if we have a closed curve and an $n$-dimensional disk in $\mathbb{R}^{n+1}$ that intersect nicely in exactly two spots, and if we have any homotopy splitting the curve and the disk without the curve passing through the boundary of the disk, then we can continuously track points in the parameter space of the curve such that the curve evaluated at these points stays arbitrarily close to the disk, and such that the parameter values eventually merge. This tracking does require us to ”slide the homotopies back and forth through time”, i.e. instead of considering $M_{t}$ and $D_{t}$, we have to consider $M_{\gamma(t)}$ and $D_{\gamma(t)}$ for some curve $\gamma$.
@@
  • 6.
    ($\left(\{1\}\times S^{1}\right)\cap Z=\emptyset$
  • 7.
    ($\left(\{0\}\times S^{1}\right)\cap Z$ is a disjoint union of two compact intervals $I_{A}$ and $I_{B}$ of positive length, each containing one point in $S$.
@@
Toward a contradiction, assume $\mathbf{A\neq B}$. By Lemma 5.33, there is a path $(\alpha,f):[0,1]\to[0,1]\times S^{1}$ from a point $(0,x)\in I_{X}^{\circ}$ to a point $(0,y)\in I_{Y}^{\circ}$ avoiding $Z^{\circ}$. By the properties of $Z$, we can assume that this path avoids $Z$. Since $Z$ is closed, we can perturb the curve $(\alpha,f)$ in the open set $\left([0,1]\times S^{1}\right)\setminus Z$ so that the curve $\alpha$ is piecewise monotone. Let $T$ be the maximum value of $\alpha$. Without loss of generality assume that $\alpha(\tfrac{1}{2})=T$, else reparametrize $(\alpha,f)$. We now define $(\alpha_{1},f_{1}),(\alpha_{2},f_{2}):[0,1]\to[0,1]\times S^{1}$ through $(\alpha_{1},f_{1})(t)\coloneqq(\alpha,f)(\tfrac{t}{2})$ and $(\alpha_{2},f_{2})\coloneqq(\alpha,f)(1-\tfrac{t}{2})$. Note that $\alpha_{1},\alpha_{2}:[0,1]\to[0,T]$ agree on endpoints and that both are piecewise monotone. We may assume that $\gamma\coloneqq\alpha_{1}=\alpha_{2}$, else we appeal to the mountain climbing theorem (5.31) for the existence of curves $\alpha^{\prime}_{1},\alpha^{\prime}_{2}:[0,1]\to[0,1]$, fixing endpoints, with $\gamma\coloneqq\alpha_{1}\circ\alpha^{\prime}_{1}=\alpha_{2}\circ\alpha^{\prime}_{2}$ and then consider $(\gamma,f_{1}\circ\alpha^{\prime}_{1})$ and $(\gamma,f_{2}\circ\alpha^{\prime}_{2})$. Let $q:\mathbb{R}\to S^{1}$ be the standard covering. Lift $f_{1}$ and $f_{2}$ to paths $\overline{f_{1}},\overline{f_{2}}:[0,1]\to\mathbb{R}$, i.e. $q\circ\overline{f_{1}}=f_{1}$ and $q\circ\overline{f_{2}}=f_{2}$. Pick the lifts so that $\overline{f_{1}}(0)<\overline{f_{2}}(0)<\overline{f_{1}}(0)+1$. Then define $h:[0,1]^{2}\to\mathbb{R}$, $h(t,x)=(1-x)\overline{f_{1}}(t)+x\overline{f_{2}}(t)$. For a given $t$, the image of $h(t,\cdot)$ corresponds to the interval between $\overline{f}_{1}(t)$ and $\overline{f}_{2}(t)$. Looking at $q(h(t,\cdot)):[0,1]\to S^{1}$ we get a continuously varying family of curves in $S^{1}$ with endpoints $f_{1}(t)$ and $f_{2}(t)$. Next we define
@@
Assume without loss of generality that $\min\left(\{t\in[0,1]\mid x_{1}(t)=x_{2}(t)\}\right)=1$. We define $\left(\tilde{M}_{t},\tilde{A}^{1}_{t},\tilde{A}^{2}_{t},\tilde{C}_{t}\right)=\tilde{H}_{t}=H_{\gamma(t)}$, and let $\tilde{D}_{t}$ be the center of mass spanning disk of $\tilde{C}_{t}$. For $t\in[0,1]$, $x_{1}(t)$ and $x_{2}(t)$ partition $S^{1}$ into two parts, say $I_{t}^{1}$ and $I_{t}^{2}$. We define $\tilde{M}_{t}^{1}$ as the concatenation of $\tilde{M}_{t}|_{I^{1}_{t}}$ and the straight line segment between $\tilde{M}_{t}(x_{1}(t))$ and $\tilde{M}_{t}(x_{2}(t))$. Analogously we define $\tilde{M}_{t}^{2}$ as the concatenation of $\tilde{M}_{t}|_{I^{2}_{t}}$ and the straight line segment between $\tilde{M}_{t}(x_{1}(t))$ and $\tilde{M}_{t}(x_{2}(t))$. These are continuously varying families of curves. Either $\tilde{M}^{1}_{1}$ or $\tilde{M}^{2}_{1}$ becomes a single point. Without loss of generality assume that $\tilde{M}^{1}_{1}$ becomes a single point. Since $\tilde{M}^{1}_{0}$ was linked to $\tilde{A}^{1}_{0}$, and since $\tilde{M}^{1}_{1}$ is a point and hence not linked to $\tilde{A}^{1}_{1}$, there must be some $t\in[0,1]$ such that $\tilde{M}^{1}_{t}$ and $\tilde{A}^{1}_{t}$ intersect. This intersection must happen on the straight line segment between $\tilde{M}_{t}(x_{1}(t))$ and $\tilde{M}_{t}(x_{2}(t))$, as $\tilde{M}_{t}$ and $\tilde{A}^{1}_{t}$ stay disjoint throughout. Let $P_{1}:=\tilde{M}_{t}(x_{1}(t))$ and $P_{2}:=\tilde{M}_{t}(x_{2}(t))$, then there is some $Q\in\tilde{A}^{1}_{t}$ on the line segment $\overline{P_{1}P_{2}}$ with $||P_{1}-Q||,||P_{2}-Q||\geq 1$. This implies $||P_{1}-P_{2}||\geq 2$. Further, we have $\mathrm{dist}(P_{1},\tilde{D}_{t})<\varepsilon^{\prime}$ and likewise $\mathrm{dist}(P_{2},\tilde{D}_{t})<\varepsilon^{\prime}$. By Proposition 4.27, we get
$\mathrm{len}(C_{\gamma(t)})=\mathrm{len}(\tilde{C}_{t})\geq 2\pi+2||P_{1}-P_{2}||-(2\pi+4)\varepsilon^{\prime}\geq 2\pi+4-\varepsilon.$

### v2

@@ Abstract
A thick link is a link in $\mathbb{R}^{3}$ such that each component of the link lies at distance at least $1$ from every other component. Strengthening the notion of thickness, we define a thickly embedded link to be a thick link whose open radius-$\tfrac{1}{2}$ normal disk bundles of all components are embedded. The Gehring ropelength problem asks how large the sum of the lengths of the components of a thick (respectively thickly embedded) link must be, given the link homotopy (respectively isotopy) class of the link. A thick homotopy (isotopy) is a link homotopy (isotopy) of a thick (thickly embedded) link that preserves thickness throughout, and such that during the homotopy the total length of the link never exceeds the initial total length. These notions of thick homotopy and isotopy are more permissive than other notions of physical link isotopies in which the length of each individual component must remain constant. We construct an explicit example of a thickly embedded 4-component link which is topologically split but cannot be split by a thick homotopy, and thick links in every homotopy class with 2 components that are non-global local minima for ropelength. This is the first time such local minima for ropelength have been explicitly constructed. In particular, we construct a thick 2-component link in the link homotopy class of the unlink which cannot be split through a thick homotopy.
@@ 1 Introduction
In 2015, Coward and Hass [5] defined the notion of a physical link isotopy. A physical link isotopy is a link isotopy that keeps a link thickly embedded throughout, as well as keeping the length of each component constant. This made it possible to study a physical knot- and link theory that corresponds to knots and links tied from real pieces of rope. In the same paper, Coward and Hass construct a $2$-component Gordian split link which they define as a thickly embedded, topologically split link that cannot be split by a physical link isotopy. More generally, a Gordian pair of links is a set of two links in $\mathbb{R}^{3}$ which are topologically equivalent, but which are not equivalent under physical deformations. Here topologically equivalent refers either to being in the same link isotopy or homotopy class, and physical deformation refers to either link homotopy or link isotopy, subject to additional geometric constraints on length and thickness. For length, there are two common constraints: either demanding that the length of each component individually does not increase throughout the homotopy/isotopy, or that their sum does not increase throughout the homotopy/isotopy. If only the sum of lengths is constrained, we say that length trading between components is permitted. The thickness constraint usually demands that links stay thick links (see Definition 4.3) or thickly embedded links (see Definition 4.4) throughout the homotopy or isotopy respectively. When considering link isotopy, sometimes additional curvature constraints are imposed.
@@
2 Main results
We briefly state the main results of this paper. All notions of thick links, thickly embedded links, thick link homotopies, thick link isotopies, and the notion of being a sink for ropelength referred to in this main results section are as in Definitions 4.3, 4.4, 4.5, and 4.14.
2.1 Result (Theorem 6.8).
There exists a $4$-component thickly embedded link that is topologically split but which cannot be split by a thick homotopy.
This proves the existence of a configuration that is physical in the strongest sense which cannot be split even under the most permissive of thickness constraints. Thus, the configuration is a Gordian split link both as a thick link and as a thickly embedded link.
2.2 Result (Theorem 7.25).
For every link homotopy class with $2$-components, there exists a thick link in that class which is a non-global, local minimum and a sink for ropelength. In particular, there are Gordian pairs in every link homotopy class with $2$ components.
This is the first time local, non-global minimizers in the Gehring ropelength problem have been proven to exist at all, and the proof is constructive. In particular this settles the question if homotopically Gordian unlinks exist in the affirmative.
2.3 Result (Theorem 7.24).
There exists a $2$-component thick link in the link homotopy class of the unlink which cannot be split by a thick link homotopy.
Together, these results demonstrate that thickness constraints meaningfully obstruct the splitting of links, and that the energy landscape of Gehring ropelength has many local minima.
@@ 3 Overview of existing results about Gordian pairs
Any Gordian pair Split link 22-component split link 22-component unlink
@@
Any Gordian pair Split link 22-component unlink All 22-component links
@@
The first Gordian unlink was constructed by Coward and Hass in 2015 [5]. In 2023, Kusner and Kusner [13] constructed a Gordian pair of thickly embedded $7$-component links. The links in the pair are link isotopic to each other, but there is no thick homotopy between them. So even though the links are equivalent in the strongest topological sense, they are not equivalent even in the weakest physical sense. It was Kusner and Kusner who first coined the term Gordian pair. In [1] Ayala constructs a thickly embedded $2$-component link in the isotopy class of the unlink which cannot be split by thick isotopy under additional curvature constraints. This was the first example of a Gordian unlink. Recently, Ayala and Hass constructed a different example of a Gordian unlink without needing any additional curvature constraints. Independently, Kusner and Kusner announced the construction of a Gordian unlink [14]. Table 1 summarizes both the existing results on Gordian pairs and the author’s contributions in this paper. The table is split into two parts, considering links with respect to isotopy and homotopy respectively. The columns of each table denote what kind of example was constructed. The rows of each table denote which geometric constraints were considered, with the strength of constraints being relaxed from top to bottom. The paper referenced in each cell is the paper in which the first example of the corresponding Gordian pair was constructed. The tables give partial orders: the further down and to the right a cell in a given table is, the harder it is to construct an example for it.
@@ 4 Preliminaries
[…]
@@
$L-L^{\prime}||<\varepsilon\implies\mathrm{len}(L^{\prime})\geq\mathrm{len}(L).$
@@
If a thick link $L$ is a minimal element of the preorder $\preceq$, then we call $L$ a sink of the ropelength functional.
@@
$(f\circ\gamma)(t)\leq f(x)\text{ for all t}\implies f\circ\gamma\text{ constant. }$
It is easy to see that every global minimum and every strict local minimum is a sink. However, even for nice spaces $X$ and continuous functions $f$, there are local minima which are not sinks and sinks which are not local minima.
@@
and with equality if and only if $\gamma(t^{\prime})\in S$. By assumption $f(\gamma(t^{\prime}))\leq c$. Therefore $f(\gamma(t^{\prime}))=c$ and hence $\gamma(t^{\prime})\in S$ . This shows that $J$ is open. Since $J$ is a non-empty, closed and open subset of $[0,1]$ we have $J=[0,1]$. So $\gamma([0,1])\subseteq S\subseteq f^{-1}(c)$, hence $f\circ\gamma$ is constant. ∎
@@ 5 Bounds on ropelength
[…]
@@ 6 An isotopically Gordian split link exists
The construction of our Gordian split link, as well as the basic structure of the proof, has a very similar flavor to that of Coward and Hass [5]. Both approaches derive a lower bound for the length needed to split a thickly embedded link by assuming that there is a thick homotopy/physical isotopy, tracking the intersections of a center-of-mass spanning disk and a curve throughout that homotopy/isotopy, and finally arguing that there needs to be some point in time where the intersections yield some lower bound through applications of propositions from Section 5. The main difference is in the book-keeping of how intersections are tracked. Coward and Hass need to resolve multiple kinds of singularities that might occur during the tracking of intersections. Our arguments use the homotopy invariance of the Brouwer degree and the mountain climbing theorem instead. We start out by stating the main proposition of this chapter, which establishes a lower-bound for the length necessary to split some thick links.
@@ 6.1 A lower bound on ropelength needed to split some thick links
The main step omitted in this sketch is how, given a homotopy $H$ preserving thickness, we can continuously keep track of the intersection points $a_{t}$ and $b_{t}$, and why we can assume that they merge eventually. In this form the claim is wrong, but we will see that it can be replaced with a more technical, true claim in the same spirit. The first tool we need to use for our book-keeping will be the homotopy invariance of the Brouwer degree of a function around 0. We will therefore briefly recall the following existence and uniqueness theorem defining the degree. This formulation is taken from Fixed Point Theory by Granas and Dugundji [10].
6.2 Theorem (Brouwer degree around 0).
@@
$\mathcal{M}=\{(U,F)\mid U\subseteq\mathbb{R}^{n}\text{ open, bounded },F:\overline{U}\to\mathbb{R}^{n}\text{ continuous},0\notin F(\partial U)\}.$
@@
We further need the following two facts about the Brouwer degree. If $F:U\to\mathbb{R}^{n}$ is $\mathcal{C}^{1}$ and if 0 is a regular value of $F$, then
@@
If $0\notin F(\partial U)$, then $0\notin F(U)\implies d(U,F)=0.$
@@
Let $D^{2}$ be a closed $2$-dimensional disk and let $K,L\subseteq D^{2}$ be closed such that $D^{2}=K\cup L$ and such that $K\cap\partial D=I_{A}\cup I_{B}$ and $L\cap\partial D=I_{X}\cup I_{Y}$ are each the disjoint union of two compact intervals of positive length, with $I_{A}\cap I_{X}$, $I_{B}\cap I_{X}$, $I_{A}\cap I_{Y}$, and $I_{B}\cap I_{Y}$ each containing exactly one element. Then either $I_{A}$ and $I_{B}$ are in the same connected component of $K$ or $I_{X}$ and $I_{Y}$ are in the same connected component of $L$.
@@
This lemma tells us that if we have a closed curve and an $n$-dimensional disk in $\mathbb{R}^{n+1}$ that intersect nicely in exactly two spots, and if we have any homotopy splitting the curve and the disk without the curve passing through the boundary of the disk, then we can continuously track points in the parameter space of the curve such that the curve evaluated at these points stays arbitrarily close to the disk, and such that the parameter values eventually merge. This tracking does require us to ’slide the homotopies back and forth through time’, i.e. instead of considering $M_{t}$ and $D_{t}$, we have to consider $M_{\gamma(t)}$ and $D_{\gamma(t)}$ for some curve $\gamma$.
@@
  • $\left(\{1\}\times S^{1}\right)\cap Z=\emptyset$
  • $\left(\{0\}\times S^{1}\right)\cap Z$ is a disjoint union of two compact intervals $I_{A}$ and $I_{B}$ of positive length, each containing one point in $S$.
@@
Toward a contradiction, assume $\mathbf{A\neq B}$. By Lemma 6.6, there is a path $(\alpha,f):[0,1]\to[0,1]\times S^{1}$ from a point $(0,x)\in I_{X}^{\circ}$ to a point $(0,y)\in I_{Y}^{\circ}$ avoiding $Z^{\circ}$. By the properties of $Z$, we can assume that this path avoids $Z$. Since $Z$ is closed, we can perturb the curve $(\alpha,f)$ in the open set $\left([0,1]\times S^{1}\right)\setminus Z$ so that the curve $\alpha$ is piecewise monotone. Let $T$ be the maximum value of $\alpha$. Without loss of generality assume that $\alpha(\tfrac{1}{2})=T$, else reparametrize $(\alpha,f)$. We now define $(\alpha_{1},f_{1}),(\alpha_{2},f_{2}):[0,1]\to[0,1]\times S^{1}$ through $(\alpha_{1},f_{1})(t)\coloneqq(\alpha,f)(\tfrac{t}{2})$ and $(\alpha_{2},f_{2})\coloneqq(\alpha,f)(1-\tfrac{t}{2})$. Note that $\alpha_{1},\alpha_{2}:[0,1]\to[0,T]$ agree on endpoints and that both are piecewise monotone. We may assume that $\gamma\coloneqq\alpha_{1}=\alpha_{2}$, else we appeal to the mountain climbing theorem (Theorem 6.4) for the existence of curves $\alpha^{\prime}_{1},\alpha^{\prime}_{2}:[0,1]\to[0,1]$, fixing endpoints, with $\gamma\coloneqq\alpha_{1}\circ\alpha^{\prime}_{1}=\alpha_{2}\circ\alpha^{\prime}_{2}$ and then consider $(\gamma,f_{1}\circ\alpha^{\prime}_{1})$ and $(\gamma,f_{2}\circ\alpha^{\prime}_{2})$. Let $q:\mathbb{R}\to S^{1}$ be the standard covering. Lift $f_{1}$ and $f_{2}$ to paths $\overline{f_{1}},\overline{f_{2}}:[0,1]\to\mathbb{R}$, i.e. $q\circ\overline{f_{1}}=f_{1}$ and $q\circ\overline{f_{2}}=f_{2}$. Pick the lifts so that $\overline{f_{1}}(0)<\overline{f_{2}}(0)<\overline{f_{1}}(0)+1$. Then define $h:[0,1]^{2}\to\mathbb{R}$, $h(t,x)=(1-x)\overline{f_{1}}(t)+x\overline{f_{2}}(t)$. For a given $t$, the image of $h(t,\cdot)$ corresponds to the interval between $\overline{f}_{1}(t)$ and $\overline{f}_{2}(t)$. Looking at $q(h(t,\cdot)):[0,1]\to S^{1}$ we get a continuously varying family of curves in $S^{1}$ with endpoints $f_{1}(t)$ and $f_{2}(t)$. Next we define
@@
[…]